# Seek-CAD: A Self-refined Generative Modeling for 3D Parametric CAD Using Local Inference via DeepSeek

### v2

@@ 3.2 Step-wise Visual Feedback with CoT for Refinement
A significant capability of reasoning LLMs lies in their ability to refine and optimize initial outputs in response to feedback. The self-refined strategy has been adopted in CAD parametric model generation [61, 2]. As CAD command sequences can be easily rendered into the perspective image of an object by using geometry kernel tools like PythonOCC, this can shift the aligned judgment from "Command-Description" to "Image-Description", which can be well handled by VLMs. In this study, we propose a novel step-wise visual feedback (SVF) strategy, which leverages not only the imagery of the object’s final form but also retains visuals depicting the intermediate configurations throughout the entire construction process. Additionally, it incorporates the CoTs from DeepSeek-R1 to guide the VLM during its evaluation.
(1) Obtain Step-wise Images.  Specifically, the initial generated CAD code $I_{0}$ is converted into a sequence ($S_{seq}$) of $n$ SSR triplets, $S_{i}$ (Recall Sec 4), as defined with $I_{0}\xrightarrow{}S_{seq}=[S_{1},S_{2},\cdot\cdot\cdot,S_{n}]$. Next, we leverage the rendering script $R(*)$ (based on PythonOCC) to subsequently compile this sequence $S$ to generate their corresponding perspective images. To better illustrate the construction process from a visual perspective, we capture both the intermediate object shapes, $M_{I}$, and the ultimate object shape, $M_{U}$. To capture $M_{I}$, for each step $k\in\{1,\dots,n\}$ in the construction sequence $S_{1}\oplus S_{2}\oplus\dots\oplus S_{k}$, the corresponding image in $M_{I}$ is rendered by $R(*)$. To maintain the correlations between entities of the object and make them visualized clearly, this rendering process would highlight the entity from the current SSR triplet $S_{k}$ while entities from all prior SSR triplets $S_{j}$ (where $j<k$) are hidden as $\bar{S_{i}}$. To achieve this, we adjust the transparency in $R_{(}*)$ to ensure that the current shape of $S_{k}$ to the $S_{j}$ intermediate shape is emphasized. To capture the ultimate shape of the object, $M_{U}$, the complete sequence $S_{seq}$ would be directly rendered without hiding any entities. The rendering process can be found in the "CAD Code Refinement" of Figure 1. Finally, each $I_{0}$ can be defined with a set of perspective images M consisting of $M_{I}$ and $M_{U}$:
@@ 4 SSR Triple Design Paradigm
(1) SSR Triple Design definition. The Sketch and Extrude (SE) paradigm persists as the predominant approach in contemporary feature-based CAD modeling [22, 57, 33, 30, 63, 49, 48, 59, 38], attributed to its versatility and facilitation of parametric editing. Current large-scale datasets for parametric sequence modeling [54, 51] also adopt this paradigm. Nevertheless, the command sets provided by preceding datasets are constrained to basic operations such as sketch and extrude, and the curves incorporated within sketches are predominantly confined to elementary types. Consequently, contemporary research endeavours leveraging these datasets tend to yield simpler and less diverse geometric forms, failing to adequately address the complexities inherent in real-world design requirements. To address this, we introduce a novel modeling paradigm called SSR (Sketch, Sketch-based feature, and Refinements), in which each modeling step is represented as a SSR triplet $S$:
@@
where $k\geq 0$, $s$ denotes a 2D sketch feature, $f\in\mathcal{F}$ is a sketch-based feature such as extrude or revolve, and $\langle r_{1},r_{2},\dots,r_{k}\rangle$ is an ordered sequence of zero or more refinement features, where each $r_{i}\in\mathcal{R}$ and $\mathcal{R}$ includes features such as fillet, chamfer, and shell. Each SSR triplet $S_{i}$ is compiled into a 3D geometry $\mathcal{S}_{i}=\texttt{CAD\_Kernel}(S_{i})$ via a CAD kernel like PythonOCC, and a complete CAD model $\mathcal{M}$, composed of $n$ SSR triples, is represented as a sequence of geometry units joined by boolean operations:
@@
Given an SSR triplet $S=(s,f,\langle r_{1},r_{2},\dots,r_{k}\rangle)$, where $s$ is the 2D sketch containing a set of primitives $\mathcal{A}=\{a_{1},a_{2},\dots,a_{m}\}$, we define the intermediate geometry $\mathcal{S}^{\prime}=\texttt{CAD\_Kernel}(S^{\prime})$, where $S^{\prime}=(s,f)$ omits the refinement operations. The resulting geometry $\mathcal{S}^{\prime}$ contains a set of primitives =$\mathcal{B}=\{b_{1},b_{2},\dots,b_{n}\}$. The CapType reference mechanism defines a mapping $\mathcal{A}\rightarrow\mathcal{B}$ as:
@@
where $C$ denotes the CapType category: START and END refer to the primitives at the start and end of the 3D operation $f$, respectively, while SWEPT refers to the primitives generated along the trajectory of the operation, as illustrated in Figure 3.
Figure 3: Illustration of the proposed CapType reference mechanism.
This mechanism allows each refinement operation $r_{i}\in\mathcal{R}$ to reference a specific primitive $b\in\mathcal{B}$ via $\phi(a,C)$, enabling precise and reliable identification. As illustrated in Figure 5 and supplementary Figure 7, the method enables the generation of complex and diverse CAD models that align well with real-world design requirements.
@@ 5 Experiments
In this section, we perform evaluations of the performance of Seek-CAD and showcase its application across a range of downstream tasks.
@@ 5.1 Experimental Settings
(1) Metrics. To evaluate how precisely the generated CAD commands depict the 3D object, we sample 2000 points separately from CAD models in Ground Truth GT and generated CAD models D. We use Chamfer Distance (CD), Hausdorff Distance (HD), and Intersection over the Ground Truth (IoGT) [2] to measure the differences between GT and D. Besides, we utilize G-Score (score from 1 to 5, allowing decimals by Gemini-2.0) to judge the alignment between the description and the perspective image of its corresponding generated CAD model. To prove that the generative model is not a mere replication of the local CAD corpus, we also compute a Novel metric as: $\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[s(I_{A},I_{B_{i}})<\tau]\geq\rho$, where $I_{A}$ is a rendering image of the generated CAD model by SeekCAD and $I_{B_{i}}$ is a rendering image of each CAD model in the local CAD corpus, and $s(*)$ is a similarity function ($\tau=0.8$, $\rho=0.8$). In practice, we use ResNet-50 [14] to encode $I_{A}$ and $I_{B_{i}}$ before calculating their similarity. Note that cases failed to compile would be excluded when calculating these metrics. Given the generated CAD commands could be failed to compile, we further add Pass@k to better understand the mechanism of components in Seek-CAD. Pass@k denotes for each description, the probability of at least one set of generated CAD commands being compiled successfully in k times generation. We report mean values under all metrics.
(2) Comparison Methods. We have witnessed two training-free efforts of generative CAD modeling with the self-refined strategy: 3D-PreMise [61] and CADCodeVerify [2]. To compare with them fairly, only the refined strategy in Seek-CAD (SVF) is separately replaced with refined strategies in 3D-PreMise and CADCodeVerify, while the rest of the components in Seek-CAD are all consistent.
@@
5.2 Experimental Results
@@
(1) Generation. To avoid data overlapping, we construct a test set consisting of an additional 500 CAD models which are completely different from 10, 000 CAD models in the local CAD corpus. Based on our framework, we compare our SVF with two other visual feedback strategies in one-step refinement, as shown in Table 1. As shown in Table 1, the SVF strategy employed in Seek-CAD significantly outperforms the feedback methods used in 3D-Premise and CADCodeVerify across all metrics. This demonstrates that the CoT and step-wise images utilized in the SVF strategy are indeed well-aligned, and the object design logic contained within the CoT empowers the VLM (Gemini-2.0) to provide clearer feedback, enabling the more precise refinement. Besides, all feedback methods based on our framework achieve close to or over 50% under the Novel metric. This indicates that while the local CAD corpus serves as a constraint on CAD command generation, it does not fully limit the creative ability of our framework to generate novel CAD models. Figure 4(a) gives comparable showcases of Ground Truth GT and the generated CAD models from Seek-CAD with different refined strategies. Compared to 3D-PreMise and CADCodeVerify, the SVF embedded in Seek-CAD makes the generated CAD model closer to GT, which again proves that SVF (Sec 3.2) is reasonable and effective to integrate step-wise images and the thought from DeepSeek-R1.
@@
(2) Refinement Step. To better understand the impact of visual feedback on CAD model generations, we explore the effects of performing 0, 1, and 2 rounds of SVF within Seek-CAD, which can be found in Table 2. Compared to Round 0, the generated CAD models show significant improvements across four metrics (CD, HD, IoGT, G-Score) after undergoing corrections in Round 1 and Round 2, which proves the effectiveness of SVF. Additionally, we found that the gains diminish significantly with an increasing round of SVF iterations. For instance, compared to Round 1, Round 2 shows only marginal improvements on the four metrics (e.g., IoGT from 0.7226 to 0.7347), but it increases the probability of code compilation failures (Pass@2 from 0.72 to 0.55). There are two main reasons for this: (i) Limited by the inherent reasoning ability of the base model (DeepSeek-R1:32B-Q4) in Seek-CAD, it is difficult for the model to make perfectly precise corrections based on SVF feedback. (ii) There is a certain probability that the generated CAD commands will fail to compile during each generation process. Each additional round of refinement would increase its probability. Hence, we set the maximum iteration step of 2 for the refinement stage to avoid excessive and potentially unnecessary modifications. Two cases are depicted in Figure 4(b). In the right-hand instance, the initially generated CAD model exhibits a reversed extrusion direction, resulting in the failure to form holes in the corners and the middle of the object, which is resolved after one round of refinement. It again demonstrates the effectiveness of the SVF strategy.
@@
(3) Ablation Study. To figure out the mechanism of Seek-CAD, we conduct ablation studies on 200 CAD models by removing some components in Seek-CAD to achieve an additional 7 frameworks tagged from A to G, which separately denote: A (w/o local CAD corpus), B (w/o knowledge constraint), C (w/o inter images in SVF), D (w/o ultimage image in SVF), E (w/o vector search in local CAD corpus), F (w/o text search in local CAD corpus), and G (w/o CoT in SVF). For the complete version of Seek-CAD, we denote it as H. Note that all comparisons in the ablation study are based on the one-time refinement. As shown in Table 3, model A completely fails to generate compilable CAD commands, indicating that within a training-free framework, the local CAD corpus indeed imposes constraints to guide model to generate CAD commands based on the SSR format. This is highly effective for the Seek-CAD model, which lacks prior knowledge of SSR, providing a technical approach for quickly achieving CAD model generations without any training. Compared to model H, the impact of models B, E, and F is primarily focused on the compilability of generated CAD commands. In particular, the substantial decline in Pass@1 for model B (from 0.68 to 0.44) highlights the critical role of the system prompt as an effective constraint during CAD command generation. Besides, models E and F also exhibit a significant decrease in four other metrics (e.g., from 0.7451 to 0.6464 in IoGT, 3.8409 to 3.6315 in G-Score). This proves conducting hybrid research in the local CAD corpus is more effective. Furthermore, we found that the image utilization strategy within SVF mainly impacts the precision of the refinement, as evidenced by model C achieving worse scores across all metrics (e.g., 0.2295 vs 0.2114 in CD) compared to model D. This validates that inter-images can provide more visual information than the ultimate-image, thereby enabling the VLM to better understand the construction process of objects and generate clear feedback to enhance further modifications. Model G shows a decline in generation quality (e.g., G-Score from 3.8409 to 3.6120) compared to model H. It indicates CoT helps Gemini-2.0 better understand the step-wise images and improve the quality of feedback. Finally, the complete model H surpasses other comparable versions across all metrics (e.g., highest score 0.7451 in IoGT), proving the design components in Seek-CAD are all effective and reasonable.
@@
(4) Various Showcases by Seek-CAD. We showcase more capabilities of our Seek-CAD,including CAD editing, Similar generations, and Gallery of generated CAD models. (i) CAD editing is an applicative sought in a wide range of industrial sectors. As Seek-CAD is a framework built on DeepSeek-R1, it supports multiple rounds of dialogue to achieve substantial editing that includes but is not limited to "Add", "Remove", and "Scale" modification. Figure 5(a) shows a demo to edit the vanilla generated CAD model based on the user’s description iteratively. (ii) Similar generations. For the same textual descriptions, Seek-CAD can generate similar CAD models as shown in Figure 5(b), which offers useful suggestions for initial CAD design. Besides, the description in Figure 5(b) contains only functional specifications of the objects, in contrast to that in Figure 5(a) which includes specific dimensional parameters. It indicates Seek-CAD is capable of interpreting functional descriptions to produce valid CAD models without relying on explicit parameter guidance. (iii) The gallery of generated CAD models. Figure 5(c) provides some generated CAD models from Seek-CAD with feeding textual descriptions from the test set. This shows that Seek-CAD is capable of generating relatively complex models, which was not demonstrated by previous methods based on SE paradigms or CadQuery.
(5) Limitations and Discussions. Seek-CAD currently has several limitations. First, the step-wise visual feedback contains biases from VLMs due to two main issues: (i) VLMs, confined to their training, struggle to accurately describe object geometries without specific-domain training, and (ii) complex models can’t be easily summarized. Second, using DeepSeek-R1 for visual feedback introduces redundant information, as full CoTs are fed to VLMs. Third, DeepSeek-R1-32B-Q4’s computational ability for geometric constraints is weaker compared to DeepSeek-R1-671B, needing more refinement for accurate CAD commands. Fourth, the CapType referencing system struggles with selecting certain topological primitives, like intersection-generated edges, which need intermediate B-Rep modeling data.
6 Conclusion
We introduce Seek-CAD, a new training-free framework for parametric CAD model generation using the open-source DeepSeek-R1 for cost-efficient inference. It refines CAD commands by integrating visual signals and CoT reasoning with Vision-Language Model evaluation. We also present a corpus of 40, 000 annotated CAD models based on the SSR design paradigm, which allows for precise operations like fillet and shell, which are absent in the traditional Sketch-Extrude paradigm datasets. Seek-CAD surpasses existing training-free methods in generation quality and geometric fidelity, pioneering the use of open-source LLMs in CAD workflows and advancing AI for design automation.

### v3

@@ 3.2 Step-wise Visual Feedback with CoT for Refinement
A significant capability of reasoning LLMs lies in their ability to refine and optimize initial outputs in response to feedback. The self-refined strategy has been adopted in CAD parametric model generation Yuan et al. (2024), Alrashedy et al. (2025). As CAD command sequences can be easily rendered into the perspective image of an object by using geometry kernel tools like PythonOCC, this can shift the aligned judgment from "Command-Description" to "Image-Description", which can be well handled by VLMs. In this study, we propose a novel step-wise visual feedback (SVF) strategy, which leverages not only the imagery of the object’s final form but also retains visuals depicting the intermediate configurations throughout the entire construction process. Additionally, it incorporates the CoTs from DeepSeek-R1 to guide the VLM during its evaluation.
(1) Obtain Step-wise Images. Specifically, the initial generated CAD code $I_{0}$ is converted into a sequence ($S_{seq}$) of $n$ SSR triplets, $S_{i}$ (Recall Sec 4), as defined with $I_{0}\xrightarrow{}S_{seq}=[S_{1},S_{2},\cdot\cdot\cdot,S_{n}]$. Next, we leverage the rendering script R()R(*) (based on PythonOCC) to subsequently compile this sequence $S$ to generate their corresponding perspective images. To better illustrate the construction process from a visual perspective, we capture both the intermediate object shapes, $M_{I}$, and the ultimate object shape, $M_{U}$. To capture $M_{I}$, for each step $k\in\{1,\dots,n\}$ in the construction sequence $S_{1}\oplus S_{2}\oplus\dots\oplus S_{k}$, the corresponding image in $M_{I}$ is rendered by R()R(*). To maintain the correlations between entities of the object and make them visualized clearly, this rendering process would highlight the entity from the current SSR triplet $S_{k}$ while entities from all prior SSR triplets $S_{j}$ (where $j<k$) are hidden as $\bar{S_{i}}$. To achieve this, we adjust the transparency in $R_{(}*)$ to ensure that the current shape of $S_{k}$ to the $S_{j}$ intermediate shape is emphasized. To capture the ultimate shape of the object, $M_{U}$, the complete sequence $S_{seq}$ would be directly rendered without hiding any entities. The rendering process can be found in the "CAD Code Refinement" of Figure 1. Finally, each $I_{0}$ can be defined with a set of perspective images M consisting of $M_{I}$ and $M_{U}$:
@@ 4 SSR Triple Design Paradigm
(1) SSR Triple Design definition. The Sketch and Extrude (SE) paradigm persists as the predominant approach in contemporary feature-based CAD modeling Khan et al. (2024a), Xu et al. (2023), Zhang et al. (2025b), Wang et al. (2024), Yavartanoo et al. (2024), attributed to its versatility and facilitation of parametric editing. Current large-scale datasets for parametric sequence modeling Wu et al. (2021), Willis et al. (2021) also adopt this paradigm. Nevertheless, the command sets provided by preceding datasets are constrained to basic operations such as sketch and extrude, and the curves incorporated within sketches are predominantly confined to elementary types. Consequently, contemporary research endeavours leveraging these datasets tend to yield simpler and less diverse geometric forms, failing to adequately address the complexities inherent in real-world design requirements. To address this, we introduce a novel modeling paradigm called SSR (Sketch, Sketch-based feature, and Refinements), in which each modeling step is represented as a SSR triplet $S$:
@@
where $k\geq 0$, $s$ denotes a 2D sketch feature, $f\in\mathcal{F}$ is a sketch-based feature such as extrude or revolve, and $\langle r_{1},r_{2},\dots,r_{k}\rangle$ is an ordered sequence of zero or more refinement features, where each $r_{i}\in\mathcal{R}$ and $\mathcal{R}$ includes features such as fillet and chamfer. Each SSR triplet $S_{i}$ is compiled into a 3D geometry $\mathcal{S}_{i}=\texttt{CAD\_Kernel}(S_{i})$ via a CAD kernel like PythonOCC, and a complete CAD model $\mathcal{M}$, composed of $n$ SSR triples, is represented as a sequence of geometry units joined by boolean operations:
@@
Figure 3: Illustration of the proposed CapType reference mechanism.
@@
Given an SSR triplet $S=(s,f,\langle r_{1},r_{2},\dots,r_{k}\rangle)$, where $s$ is the 2D sketch containing a set of primitives $\mathcal{A}=\{a_{1},a_{2},\dots,a_{m}\}$, we define the intermediate geometry $\mathcal{S}^{\prime}=\texttt{CAD\_Kernel}(S^{\prime})$, where $S^{\prime}=(s,f)$ omits the refinement operations. The resulting geometry $\mathcal{S}^{\prime}$ contains a set of primitives $\mathcal{B}=\{b_{1},b_{2},\dots,b_{n}\}$. The CapType reference mechanism defines a mapping $\mathcal{A}\rightarrow\mathcal{B}$ as:
@@
where $C$ denotes the CapType category: START and END refer to the primitives at the start and end of the 3D operation $f$, respectively, while SWEPT refers to the primitives generated along the trajectory of the operation, as illustrated in Figure 3. This mechanism allows each refinement operation $r_{i}\in\mathcal{R}$ to reference a specific primitive $b\in\mathcal{B}$ via $\phi(a,C)$, enabling precise and reliable identification. As illustrated in Figure 5(c) and Figure 7, the method enables the generation of complex and diverse CAD models that align well with real-world design requirements.
@@ 5.1 Experimental Settings
(1) Metrics. To evaluate how precisely the generated CAD commands depict the 3D object, we sample 2000 points separately from CAD models in Ground Truth GT and generated CAD models D. We use Chamfer Distance (CD), Hausdorff Distance (HD), and Intersection over the Ground Truth (IoGT) Alrashedy et al. (2025) to measure the differences between GT and D. Besides, we utilize G-Score (score from 1 to 5, allowing decimals by Gemini-2.0) to judge the alignment between the description and the perspective image of its corresponding generated CAD model. To prove that the generative model is not a mere replication of the local CAD corpus, we also compute a Novel metric as: $\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[s(I_{A},I_{B_{i}})<\tau]\geq\rho$, where $I_{A}$ is a rendering image of the generated CAD model by SeekCAD and $I_{B_{i}}$ is a rendering image of each CAD model in the local CAD corpus, and s()s(*) is a similarity function ($\tau=0.8$, $\rho=0.8$). In practice, we use ResNet-50 He et al. (2016) to encode $I_{A}$ and $I_{B_{i}}$ before calculating their similarity. Note that cases failed to compile would be excluded when calculating these metrics. Given the generated CAD commands could be failed to compile, we further add Pass@k to better understand the mechanism of components in Seek-CAD. Pass@k denotes for each description, the probability of at least one set of generated CAD commands being compiled successfully in k times generation. We report mean values under all metrics.
(2) Comparison Methods. We have witnessed two training-free efforts of generative CAD modeling with the self-refined strategy: 3D-PreMise Yuan et al. (2024) and CADCodeVerify Alrashedy et al. (2025). To compare with them fairly, only the refined strategy in Seek-CAD (SVF) is separately replaced with refined strategies in 3D-PreMise and CADCodeVerify, while the rest of the components in Seek-CAD are all consistent. Besides, we choose the latest finetune strategy (CAD-Llama Li et al. (2025b)) as a competitor to show the capability of Seek-CAD.
(3) Implementation Details. We deploy DeepSeek-R1:32B in Q4 quantization version as a backbone of Seek-CAD on one NVIDIA RTX 3090 GPU with Ollama Ollama . The context length is set to 15,000 to endow inference speed of 21.78 tokens per second. For the inference stage, we set temperature $T=0.7$, top-p as 0.8 to make it capable of generating diverse CAD models in different trials. For RAG settings, we employ Dify Dify and bge-m3 Chen et al. (2024), details refer to Appendix A.4. For the refinement stage, we adopt the Gemini-2.0 API Team et al. (2024) from Google AI to judge the discrepancies between step-wise perspective images and their corresponding thoughts (DeepSeek-R1:32B-Q4). The feedback from Gemini would be delivered back to Seek-CAD to again refine the generated CAD commands. We set the maximize iteration step of 1 for the refinement stage.
5.2 Experimental Results
(1) Generation. To avoid data overlapping, we construct a test set consisting of an additional 500 CAD models which are completely different from 10, 000 CAD models in the local CAD corpus.
Table 1: The quantitative results of generation ability tested on 500 CAD models.
@@ 5.2 Experimental Results
Finetune CAD-Llama 0.2147 0.5864 0.7023 3.3385 77.64%
Training-free 3D-PreMise 0.2203 0.6137 0.6315 3.2022 49.57%
CADCodeVerify 0.2164 0.5917 0.6562 3.3927 55.38%
Seek-CAD (Ours) 0.1979 0.5566 0.7226 3.5185 64.04%
Based on our framework, we compare our SVF with two other visual feedback strategies in one-step refinement. As shown in Table 1, the SVF strategy employed in Seek-CAD significantly outperforms the feedback methods used in 3D-Premise and CADCodeVerify across all metrics. This demonstrates that the CoT and step-wise images utilized in the SVF strategy are indeed well-aligned, and the object design logic contained within the CoT empowers the VLM (Gemini-2.0) to provide clearer feedback, enabling the more precise refinement. Besides, all feedback methods based on our framework achieve close to or over 50% under the Novel metric. This indicates that while the local CAD corpus serves as a constraint on CAD command generation, it does not fully limit the creative ability of our framework to generate novel CAD models. Figure 4(a) gives comparable showcases of Ground Truth GT and the generated CAD models from Seek-CAD with different refined strategies. Compared to 3D-PreMise and CADCodeVerify, the SVF embedded in Seek-CAD makes the generated CAD model closer to GT, which again proves that SVF (Sec 3.2) is reasonable and effective to integrate step-wise images and the thought from DeepSeek-R1. Besides, Seek-CAD shows superior performance across CD, HD, IoGT, and G-Score metrics compared to CAD-Llama. This highlights its ability to prioritize geometric accuracy and achieve higher text-matching fidelity, proving a stronger semantic understanding. Although it scores lower on novelty than CAD-Llama, this trade-off results in models that are geometrically closer to the ground truth (Figure 4(a)). Crucially, Seek-CAD’s ability to completely bypass the training phase shows it can rival the overall performance of fine-tuned models, presenting a significant advantage when training resources are limited.
@@
Table 2: The quantitative results of comparing refinement rounds tested on 500 CAD models.
[…]
0 0.77 0.2275 0.6194 0.6183 3.1401
1 0.72 0.1979 0.5566 0.7226 3.5185
2 0.55 0.1966 0.5548 0.7347 3.5314
@@
(2) Refinement Step. To better understand the impact of visual feedback on CAD model generations, we explore the effects of performing 0, 1, and 2 rounds of SVF within Seek-CAD, which can be found in Table 2. Compared to Round 0, the generated CAD models show significant improvements across four metrics (CD, HD, IoGT, G-Score) after undergoing corrections in Round 1 and Round 2, which proves the effectiveness of SVF. Additionally, we found that the gains diminish significantly with an increasing round of SVF iterations. For instance, compared to Round 1, Round 2 shows only marginal improvements on the four metrics (e.g., IoGT from 0.7226 to 0.7347), but it increases the probability of code compilation failures (Pass@2 from 0.72 to 0.55). There are two main reasons for this: (i) Limited by the inherent reasoning ability of the base model (DeepSeek-R1:32B-Q4) in Seek-CAD, it is difficult for the model to make perfectly precise corrections based on SVF feedback. (ii) There is a certain probability that the generated CAD commands will fail to compile during each generation process. Each additional round of refinement would increase its probability. Hence, we set the maximum iteration step of 1 for the refinement stage to avoid excessive and potentially unnecessary modifications. Two cases are depicted in Figure 4(b). In the right-hand instance, the initially generated CAD model exhibits a reversed extrusion direction, resulting in the failure to form holes in the corners and the middle of the object, which is resolved after one round of refinement. It again demonstrates the effectiveness of the SVF strategy.
(3) Ablation Study. To figure out the mechanism of Seek-CAD, we conduct ablation studies on 200 CAD models by removing some components in Seek-CAD to achieve an additional 7 frameworks tagged from A to G, which separately denote: A (w/o local CAD corpus), B (w/o knowledge constraint), C (w/o inter images in SVF), D (w/o ultimage image in SVF), E (w/o vector search in local CAD corpus), F (w/o text search in local CAD corpus), and G (w/o CoT in SVF). For the complete version of Seek-CAD, we denote it as H. Note that all comparisons in the ablation study are based on the one-time refinement. As shown in Table 3, model A completely fails to generate compilable CAD commands, indicating that within a training-free framework, the local CAD corpus indeed imposes constraints to guide model to generate CAD commands based on the SSR format. This is highly effective for the Seek-CAD model, which lacks prior knowledge of SSR, providing a technical approach for quickly achieving CAD model generations without any training. Compared to model H, the impact of models B, E, and F is primarily focused on the compilability of generated CAD commands. In particular, the substantial decline in Pass@1 for model B (from 0.68 to 0.44) highlights the critical role of the system prompt as an effective constraint during CAD command generation. Besides, models E and F also exhibit a significant decrease in four other metrics (e.g., from 0.7451 to 0.6464 in IoGT, 3.8409 to 3.6315 in G-Score).
@@
This proves conducting hybrid research in the local CAD corpus is more effective. Furthermore, we found that the image utilization strategy within SVF mainly impacts the precision of the refinement, as evidenced by model C achieving worse scores across all metrics (e.g., 0.2295 vs 0.2114 in CD) compared to model D. This validates that inter-images can provide more visual information than the ultimate-image, thereby enabling the VLM to better understand the construction process of objects and generate clear feedback to enhance further modifications. Model G shows a decline in generation quality (e.g., G-Score from 3.8409 to 3.6120) compared to model H. It indicates CoT helps Gemini-2.0 better understand the step-wise images and improve the quality of feedback. Finally, the complete model H surpasses other comparable versions across all metrics (e.g., highest score 0.7451 in IoGT), proving the design components in Seek-CAD are all effective and reasonable.
(4) VLM feedback quality. To provide a deeper insight, we conducted a new statistical analysis of the VLM feedback, categorizing the responses into three types: "Yes" (Aligned),"No" (Misaligned), and"Unsure" .We performed this analysis on the 500 CAD samples from our test set (utilizing the optimal refinement step $N=1$). The detailed statistics are presented in Table 4.
Table 4: The quantitative analysis of the VLM feedback on 500 CAD models.
Helpful Useless Harmful
("Yes"/"No") "Unsure" -
(67/374) 59 -
88.2% 11.8% -
Based on it, the impact of each type can be summarized: (i) "Unsure" (Useless): This category represents cases where the VLM failed to make a definitive judgment. In these cases, the feedback is useless (ineffective), as it does not trigger a meaningful correction, but it effectively acts as a "pass" and does not actively mislead the model. (ii) "Yes" and "No" (Helpful): We carefully verified samples from these two categories by human. We found that these judgments were overwhelmingly valid. "Yes" correctly validated accurate models, and "No" indeed identified logical discrepancies between the CoT and the images. (iii) "Harmful": We acknowledge that harmful hallucinations (e.g., claiming a curve is "not smooth" when it is) are theoretically possible and might appear in larger-scale testing, our manual verification indicates they were not a dominant factor here.
(5) Performance vary with CAD model complexity. To further uncover the capabilities of Seek- CAD, we use the total number of CAD commands within the SSR paradigm as the metric for complexity. Specifically, we categorized the generated CAD models into three groups based on their command sequence length: Low [0, 30], Medium [31, 70], and High [71, ). The results can be found in Table 5.
Table 5: The results of performance vary with CAD model complexity.
[…]
$[0,30]$ 0.1898 0.5131 0.7356 3.9324 56.25%
$[31,70]$ 0.2001 0.5704 0.7021 3.4935 61.76%
$[71,)$ 0.2093 0.5924 0.6759 3.1133 69.34%
The increased sequence length makes the generation task harder, increasing the difficulty of perfectly matching the target Ground Truth (leading to lower geometric scores). However, it also introduces more degrees of freedom in the generation process. This enhances the diversity of the generated models, thereby boosting their Novelty.
(6) Robustness. We conduct a new set of comparative experiments conducted on the DeepCAD dataset Wu et al. (2021) to validate the robustness of our framework. For this experiment, wereplaced our local RAG corpus with 4,000 CAD models collected from the DeepCAD datasetand another 300 CAD models (different from 4,000 CAD models in the local RAG corpus) as the test set. The comparative results can be found in Table 6. It demonstrates that even on the simpler SE-paradigm dataset, Seek-CAD continues to outperform the baselines across evaluation metrics. This validates the feasibility and robustness of our framework, confirming that its effectiveness is not solely dependent on our proposed SSR-based dataset.
(7) Various Showcases by Seek-CAD. We showcase more capabilities of our Seek-CAD,including CAD editing, Similar generations, and Gallery of generated CAD models. (i) CAD editing is an applicative sought in a wide range of industrial sectors. As Seek-CAD is a framework built on DeepSeek-R1, it supports multiple rounds of dialogue to achieve substantial editing that includes but is not limited to "Add", "Remove", and "Scale" modification.
Table 6: The quantitative results of generation ability tested on the DeepCAD dataset.
[…]
3D-PreMise 0.2079 0.5823 0.7375 3.3192 48.89%
CADCodeVerify 0.2035 0.5745 0.7531 3.6452 51.18%
Seek-CAD(ours) 0.1811 0.5231 0.8095 4.0604 54.78%
Figure 5(a) shows a demo to edit the vanilla generated CAD model based on the user’s description iteratively. (ii) Similar generations. For the same textual descriptions, Seek-CAD can generate similar CAD models as shown in Figure 5(b), which offers useful suggestions for initial CAD design. Besides, the description in Figure 5(b) contains only functional specifications of the objects, in contrast to that in Figure 5(a) which includes specific dimensional parameters. It indicates Seek-CAD is capable of interpreting functional descriptions to produce valid CAD models without relying on explicit parameter guidance. (iii) The gallery of generated CAD models. Figure 5(c) provides some generated CAD models from Seek-CAD with feeding textual descriptions from the test set. This shows that Seek-CAD is capable of generating relatively complex models, which was not demonstrated by previous methods based on SE paradigms or CadQuery.
@@
(8) Limitations and Discussions. Seek-CAD has several limitations. First, the step-wise visual feedback contains biases from VLMs due to two main issues: (i) VLMs, confined to their training, struggle to accurately describe object geometries without specific-domain training, and (ii) complex models can’t be easily summarized. Second, using DeepSeek-R1 for visual feedback introduces redundant information, as full CoTs are fed to VLMs. Third, DeepSeek-R1-32B-Q4 exhibits weaker computational capabilities for geometric constraints compared to DeepSeek-R1-671B model. Specifically, even when the VLM provides correct feedback, the 32B-Q4 model sometimes struggles to generate the precise parameters required to satisfy these constraints. This limitation could be significantly mitigated when scaling up to DeepSeek-R1-671B. Fourth, the CapType system struggles with selecting certain topological primitives, like intersection-generated edges, which need intermediate B-Rep modeling data. Specifically, if a logical error is related to a primitive that CapType cannot reference (e.g., "the fillet on that intersection is missing"), the loop cannot currently fix it because the system has no mechanism to point to that specific location.